OpenStudy (jadebirdfly):

Please help I will fan and medal!!! 0.30+0.31+0.32+0.33+0.34+....0.47+0.48+0.49+0.50

OpenStudy (phi):

you could write each term as 0.30+ x where x is 0.00 0.01, 0.02, etc you have 21 terms so you have 0.30*21 + 1/100 * sum 0,1,2,3,... 20 the what is the sum of the integers from 1 to 20 ? (there is a formula for this) ?

OpenStudy (jadebirdfly):

its ok i figured it out from there

OpenStudy (phi):

for example, if we do just the first 3 terms, we would write it like this 0.3 + 0 + 0.3 + 0.01 + 0.3 + 0.02 when adding we can reorder , right ? so we could write this as 0.3 + 0.3 + 0.3 + 0 + 0.01 + 0.02 we have add 0.3 three times. in other words we can write that as 0.3 * 3 + 0 + 0.01 +0.02 now look at 0+0.01+0.02 we could (to make it look like whole numbers) multiply and divide each by 100: \[0.00 \cdot \frac{100}{100} +0 .01 \cdot \frac{100}{100} +0.02\cdot \frac{100}{100} \] which is the same as \[ \frac{0}{100} + \frac{1}{100} +\frac{2}{100} \\ \frac{1}{100} \left( 0+1+2\right)\] and we have \[ 0.3 \cdot 3 + \frac{1}{100} \sum_{n=0}^3n \]